P=100-2Q where Q is total quantity demanded for both firms 1 and 2, respectively. The firms 1 marginal cost is given by MC1(Q1)=2Q1. The firms 2 marginal cost is given by MC2(Q2)=4Q2. Based on this information firm 1 and firm 2's reaction functions are?
Answer #1

Each firm tries to maximize profits based on the condition MR = MC

Firm 1:

TR = PQ1 = (100 - 2Q1 - 2Q2)*Q1

MR = dTR/dQ1 = 100 - 4Q1 - 2Q2

MC = 2Q1

Thus, at profit max we have, 100 - 4Q1 - 2Q2 = 2Q1

=> 6Q1 = 100 - 2Q2

=> Q1 = (50 - Q2)/3

Firm 2:

TR = PQ2 = (100 - 2Q1 - 2Q2)*Q2

MR = dTR/dQ2 = 100 - 4Q2 - 2Q1

MC = 4Q2

Thus, at profit max we have, 100 - 4Q2 - 2Q1 = 4Q2

=> 8Q2 = 100 - 2Q1

=> Q2 = (50 - Q1)/4